OpenStudy (anonymous):

angle difference identity to find the exact value of; cos 75°.

OpenStudy (igreen):

Can't you just use a calculator to find \(\sf cos(75^o)\)?

OpenStudy (anonymous):

i need to use angle difference identity to solve it.

OpenStudy (anonymous):

yes look cos(120-45)=cos(120)cos(45)+sin(120)sin(45)

OpenStudy (anonymous):

is that what you want

OpenStudy (anonymous):

it ca be written as cos(30+45)

OpenStudy (igreen):

Yes..^^^

OpenStudy (igreen):

We have: \(\sf cos(30^o + 45^o) = cos(30^o)cos(45^o) - sin(30^o)sin(45^o)\) Which gives us: \(\dfrac{\sqrt3}{2} \times \dfrac{\sqrt2}{2} - \dfrac{1}{2} \times \dfrac{\sqrt2}{2}\)

OpenStudy (igreen):

Can you simplify that? @secretslocked

OpenStudy (anonymous):

igreen did you read ? angle "DIFFERENCE"

OpenStudy (igreen):

Oh..I'm doing Angle-Sum..sorry.

OpenStudy (anonymous):

so you should do cos(120-45)=-cos(60)cos(45)+sin(60)sin(45)

OpenStudy (anonymous):

@joyraheb why 60?

OpenStudy (anonymous):

because cos(120)=cos(180-60)=-cos(60) understood?

OpenStudy (anonymous):

sort of

OpenStudy (anonymous):

and sin(120)=sin(180-60)=sin(60)

OpenStudy (anonymous):

and you know sin(60)=√3/2 and cos(60)=1/2

OpenStudy (anonymous):

and cos(45)=sin(45)=√2/2 and thats it now just replace and youll get cos(75)